OpenStudy (anonymous):

Factor completely: 4x3 + 8x2 - 25x - 50

OpenStudy (anonymous):

Damn im sorry wait Factor completely: 4x^3 + 8x^2 - 25x - 50 i just looked at it

OpenStudy (anonymous):

\[4x^3 + 8x^2 - 25x - 50=4x^2(X + 2) - 25(x +2) =(4x^2-25)(X + 2) = (2X+5)(2X-5)(X+2)\] THUS (2X+5)(2X-5)(X+2) ARE THE REQ FACTORS.